If tan x + cot x is equal to 2 then find the value of tan^2 x - cot^2 x  is???
 

Grade:12

1 Answers

Aditya Gupta
5 years ago
hello nitin, let tanx=y
so, y+1/y=2
on solving y^2+1=2y
or, y=1
so, tan^2 x - cot^2 x=y^2-1/y^2
=1-1
= zero (0)